# supplementary

### v1

@@ 2.2. Spatial and Non-Spatial Bayesian Analysis Models
$\begin{split}Y&=X\beta_{S}+\gamma+\nu_{S},\quad\gamma=\Phi M,\quad\nu_{S}=\Psi M,\quad\Phi=(\phi_{1},\ldots,\phi_{k}),\quad\Psi=(\psi_{1},\ldots,\psi_{k}),\\ \phi_{i}&\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}N_{n}(0_{n},r_{i}V_{\phi}),\quad\psi_{i}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}N_{n}(0_{n},(1-r_{i})I_{n})\;\text{ for $i=1,\ldots,k$,}\end{split}$ (6)
@@
where $\gamma$ is $n\times k$ comprising spatial effects, $M$ is $k\times k$ and invertible that induces correlation between outcomes, $W_{\phi}$ is known $n\times n$ and positive definite capturing spatial autocorrelation, and $R=\mbox{diag}(r_{1},\ldots,r_{k})$ with $r_{i}\in(0,1)$ for $i=1,\ldots k$. The analysis parameters $\{\beta_{S},M,R\}$ mimic the role of $\{\beta,A,P\}$ in the data-generation model, while the spatial structure of $W_{\phi}$ parallels $V_{\phi}$. Latent spatial structure is introduced by constructing $W_{\phi}^{-1}$ using, for example, the conditionally autoregressive model (CAR), simultaneous autoregressive model (SAR), or the directed acyclic graph auto-regressive model (DAGAR) developed in Datta et al. (2019).
[…]
@@
$\begin{split}\pi(\beta_{S},\gamma,M,R)&\propto\mbox{N}_{nk}(\operatorname{vec}(\gamma)\,|\,0,M^{\mathrm{\scriptscriptstyle T}}RM\otimes W_{\phi})\\ &\qquad\times\mbox{IW}(M^{\mathrm{\scriptscriptstyle T}}M\,|\,v,v\Sigma_{0})\times\mbox{PC}(R\,|\,\lambda_{R})\times\Pi_{i=1}^{k}M_{ii}^{k-i+1}.\end{split}$ (7)
For tractable analysis, we consider the posterior mean of $\beta_{S}$ conditioned on $M$ and $R$ as a point estimator for $\beta$. Section 3.1 characterizes spatial confounding in the spatial analysis model through the posterior variance of $\beta_{S}$ and bias of the conditioned spatial estimator for fixed $(X_{1},Z)$, similar to the original analysis of Hodges and Reich (2010). Section 3.2 shows that in the presence of a spatial confounder $Z$, the conditioned spatial estimator nevertheless retains superior estimation performance compared to the non-spatial estimator under the data generation model in (2).
@@ 3. Multivariate Spatial Confounding
3.1. Spatial Confounding in the Analysis Model
Under the non-spatial analysis model in (4), a natural point estimate for $\beta$ is $\widehat{\beta}_{NS}=\mathbb{E}[\beta_{NS}\,|\,Y,X_{1}]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}Y$, which is equivalent to separate univariate ordinary least squares estimates for each outcome. Since $\widehat{\beta}_{NS}$ does not depend on $\Sigma_{NS}$, the dependence between the outcomes does not have an effect on the point estimate of $\beta$ and, subsequently, the bias introduced by the confounder $Z$. Under the data generation model in (2),
$\operatorname{Bias}(\widehat{\beta}_{NS}\,|\,X_{1},Z)=\mathbb{E}[\widehat{\beta}_{NS}-\beta\,|\,X_{1},Z]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}Z.\$ (8)
For the spatial analysis model given by (6), let $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $B_{1}=I_{k}\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $B_{2}^{-1}=M^{-1}(I_{k}-R)^{-1}M^{-\mathrm{\scriptscriptstyle T}}\otimes(I-H)+M^{-1}R^{-1}M^{-\mathrm{\scriptscriptstyle T}}\otimes W_{\phi}^{-1}$, and $\widehat{\gamma}=\mathbb{E}[\gamma\,|\,Y,M,R,X_{1}]$. Derivation of the conditional posterior $\pi(\beta_{S},\gamma\,|\,Y,M,R,X_{1})$ in Appendix A reveals that $\operatorname{vec}(\beta_{S})\,|\,Y,M,R,X_{1}\sim\mbox{N}_{pk}(\operatorname{vec}(\widetilde{\beta}_{S}),M^{\mathrm{\scriptscriptstyle T}}(I_{k}-R)M\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+B_{1}B_{2}B^{\mathrm{\scriptscriptstyle T}}_{1})$ where the conditional spatial estimator $\widetilde{\beta}_{S}=\mathbb{E}[\beta_{S}\,|\,Y,M,R,X_{1}]$ is
$\widetilde{\beta}_{S}=\widehat{\beta}_{NS}-(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\widehat{\gamma},\quad\operatorname{vec}({\widehat{\gamma}})=(M^{\mathrm{\scriptscriptstyle T}}\otimes I_{n})B^{-1}\operatorname{vec}\left((I_{n}-H)YM^{-1}\right),$ (9)
and $B=I_{k}\otimes(I_{n}-H)+R^{-1}(I_{k}-R)\otimes W_{\phi}^{-1}$. The term $(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\widehat{\gamma}$ illuminates the exact contribution of adding a spatial effect to the estimation of $\beta$ that is derived from an inseparable interaction of latent spatial autocorrelation and dependence between results through $W_{\phi}^{-1}$, $M$, and $R$. The spatial effects $\gamma$ absorb spatial information from $Y$ through the residuals $(I_{n}-H)Y$, as $\widetilde{\beta}_{S}$ can be interpreted as the estimate of $\beta$ under the non-spatial model in (4) with $Y_{\star}=Y-\mathbb{E}[\gamma\,|\,Y,M,R,X_{1}]$ as the response instead of $Y$. Similar to how restricted spatial regression preserves the non-spatial estimate of $\beta$ under a spatial model, the non-spatial estimator can be recovered under the spatial analysis model in (6) by the relation $\widehat{\beta}_{NS}=\mathbb{E}[\beta_{S}+(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\gamma\,|\,Y,M,R,X_{1}]$.
An immediate consequence of (9) is that the bias of $\widetilde{\beta}_{S}$ under (2) can be decomposed as
$\begin{split}\operatorname{vec}(\operatorname{Bias}(\widetilde{\beta}_{S}\,|\,X_{1},Z))&=\operatorname{vec}(\operatorname{Bias}(\widehat{\beta}_{NS}\,|\,X_{1},Z))\\ &\qquad\qquad-(M^{\mathrm{\scriptscriptstyle T}}\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}})B^{-1}\operatorname{vec}((I_{n}-H)ZM^{-1}),\end{split}$ (10)
where the second term is the additional bias induced by the addition of the spatial effects $\gamma$. If the columns of $Z$ are linear combinations of the columns of $X$, then $(I-H)Z=0$ and $\operatorname{Bias}(\widetilde{\beta}_{S})=\operatorname{Bias}(\widehat{\beta}_{NS})$. Bias already exists in the non-spatial model from collinearity between $X$ and $Z$, and the spatial random effect adjusts the bias based on a product between $X$ and the projection of $Z$ onto the orthogonal complement of $X$ that is weighted according to both spatial structure and between-outcome covariance. Whether or not the adjustment exacerbates or counterbalances the non-spatial bias depends on the exact structure of $X$ and $Z$. The non-spatial estimator $\widehat{\beta}_{NS}$ is unbiased if and only if $X^{\mathrm{\scriptscriptstyle T}}Z=0$, while $\widetilde{\beta}_{S}$ is unbiased if and only if $\operatorname{vec}(X^{\mathrm{\scriptscriptstyle T}}Z)=(M^{\mathrm{\scriptscriptstyle T}}\otimes X^{\mathrm{\scriptscriptstyle T}})B^{-1}\operatorname{vec}((I-H)ZM^{-1})$.
Hodges and Reich (2010) argue that severe variance inflation from spatial confounding occurs when the ratio of non-spatial to spatial variance is near 0. We show that the inflation of $\operatorname{Var}(\operatorname{vec}(\beta_{S})\,|\,Y,M,R,X_{1})$ in the analysis model given by (6) is limited when $W_{\phi}^{-1}$ is positive definite. By spectral decomposition, there exist an orthogonal matrix $Q\in\mathbb{R}^{n\times n}$ and a diagonal matrix $D=\mbox{diag}(d_{1},\ldots,d_{n})$ with $d_{i}\geq 0$ for all $i=1,\ldots,n$ such that $W_{\phi}^{1/2}(I-H)W_{\phi}^{1/2}=QDQ^{\mathrm{\scriptscriptstyle T}}$. The derivations in Appendix A reveal that $\operatorname{Var}(\operatorname{vec}(\beta_{S})\,|\,Y,M,R,X_{1})=M^{\mathrm{\scriptscriptstyle T}}(I_{k}-R)M\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+S\left((I_{k}-R)^{-1}\otimes D+R^{-1}\otimes I_{n}\right)^{-1}S^{\mathrm{\scriptscriptstyle T}}$, where $U=W_{\phi}^{1/2}Q$ and $S=M^{\mathrm{\scriptscriptstyle T}}\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}U$. Let $R=rI_{k}$ such that $r\in(0,1)$ represents the proportion of spatial variation. Spatial smoothing increases as $r\rightarrow 1^{-}$, but for any $d_{i}$, $\lim_{r\rightarrow 1^{-}}\frac{1}{1/r+d_{i}/(1-r)}=\mbox{I}(d_{i}=0)$, where $\mbox{I}(\cdot)$ is the indicator function. Let $D^{\star}$ be a $n\times n$ diagonal matrix such that $D^{\star}_{ii}=\mbox{I}(d_{i}>0)$ for $i=1,\ldots,n$. Then, $\lim_{r\rightarrow 1^{-}}\left((I_{k}-R)^{-1}\otimes D+R^{-1}\otimes I_{n}\right)^{-1}=I_{k}\otimes(I_{n}-D^{\star})$ and the posterior conditional variance approaches the limit
$\lim_{r\rightarrow 1^{-}}\operatorname{Var}(\operatorname{vec}(\beta_{S})\,|\,Y,M,r,X_{1})=M^{\mathrm{\scriptscriptstyle T}}M\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}(W_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}})X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1},$ (11)
suggesting that setting $W_{\phi}^{-1}$ as positive definite guards against extreme variance inflation from spatial confounding. Furthermore, $\lim_{r\rightarrow 1^{-}}\widetilde{\beta}_{S}=\widehat{\beta}_{NS}-(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}UD^{\star}U^{-1}Y$ from similar calculations. In a univariate BYM2 model, Wu and Banerjee (2025) provides the limiting distribution of the total error variance as the smoothing ratio approaches zero. In our setting, the limiting distribution of $M^{\mathrm{\scriptscriptstyle T}}M$, derived in Appendix B as
$\lim_{r\rightarrow 1^{-}}\pi(M^{\mathrm{\scriptscriptstyle T}}M\,|\,Y,r,X_{1})=\mbox{IW}(M^{\mathrm{\scriptscriptstyle T}}M\,|\,v+n-p-1,v\Sigma_{0}+Y^{\mathrm{\scriptscriptstyle T}}U^{-\mathrm{\scriptscriptstyle T}}D^{\star}U^{-1}Y),$ (12)
@@ 3.1. Spatial Confounding in the Analysis Model
illuminates the limiting effect of maximal spatial smoothing on the scale of the conditional variance and between-outcome dependence. The asymptotic results of (11) and (12) indicate that the posterior variance of $\beta_{S}$ does not explode when the noise-to-signal ratio is near zero, contradicting the argument of Hodges and Reich (2010) that variance inflation from spatial confounding is comparable to multicollinearity in linear models. Statistical inference remains intact under extreme spatial smoothing, unlike the pathological case of perfect multicollinearity in covariates. Analysts should assess the degree of spatial smoothing present in the data to assess whether the spatial effect is problematic for the spatial analysis model.
Multiple authors point out that restricted spatial regression can lead to misleading statistical inference. For example, Khan and Calder (2022) shows that the marginal posterior variance in restricted spatial regression is less than the posterior variance in a non-spatial model, contradicting a central hypothesis of restricted spatial regression models. From a frequentist data-generation perspective, Zimmerman and Ver Hoef (2022) points out that conditional on the variance parameter, restricted spatial regression underestimates the variability of the fixed effects’ estimator in the true data generation model. In the following section, we reinforce the notion that restricted spatial regression methods are unnecessary by showing that under the data generation model given by (2), the Bayesian spatial estimator $\widetilde{\beta}_{S}$ is more efficient than the non-spatial estimator $\widehat{\beta}_{NS}$. We connect the analysis model and the data generation perspectives by demonstrating that the posterior variance in the analysis model captures the estimator’s variance in the data generation model.
3.2. Statistical Inference Over Multiple Datasets
Although predictors are often considered as fixed variables when evaluating estimation of the fixed effects $\beta$, treating $X_{1}$ as stochastic is also logical when covariates, such as population characteristics, are measured or estimated with finite precision. Hence, we consider the bias and mean-squared error of point estimates of $\beta$ over repeated realizations of $(X_{1},Z,Y)$ to account for the variance in the predictors and confounding effects. If the data generation model satisfies $\mathbb{E}[Z\,|\,X_{1}]=X\delta$, where $\delta$ is any $(p+1)\times k$ matrix, then (10) implies $\mathbb{E}[\widetilde{\beta}_{S}]=\mathbb{E}[\widehat{\beta}_{NS}]=\beta+\delta$, where the expectation is taken over the joint distribution of $(X_{1},Y,Z)$. In this case, $\widetilde{\beta}_{S}$ and $\widehat{\beta}_{NS}$ are equally biased regardless of the values of $M$, $R$, and $W_{\phi}^{-1}$ used to construct $\widetilde{\beta}_{S}$. Analysis in Appendix C reveals that if $\operatorname{Var}(\operatorname{vec}(Y)\,|\,X_{1})=A^{\mathrm{\scriptscriptstyle T}}PA\otimes V_{\phi}+A^{\mathrm{\scriptscriptstyle T}}(I_{k}-P)A\otimes I_{n}$, $V_{\phi}=W_{\phi}$, and $\widetilde{\beta}_{S}=\mathbb{E}[\beta_{S}\,|\,Y,M=A,R=P,X_{1}]$, then $\operatorname{Var}(\widetilde{\beta}_{S,ij}\,|\,X_{1})<\operatorname{Var}(\widehat{\beta}_{NS,ij}\,|\,X_{1})$, and if $\mathbb{E}[Z\,|\,X_{1}]=X\delta$ as well,
$\operatorname{Var}(\widetilde{\beta}_{S,ij})<\operatorname{Var}(\widehat{\beta}_{NS,ij})\;\text{and}\;\mathbb{E}\left[(\widetilde{\beta}_{S,ij}-\beta_{ij})^{2}\right]<\mathbb{E}\left[(\widehat{\beta}_{NS,ij}-\beta_{ij})^{2}\right]$ (13)

### v2

@@ 2.3.2. Spatial Bayesian Analysis
$\begin{split}Y&=XB_{S}+G+E_{S},\quad G=\Phi M,\quad E_{S}=\Psi M,\quad\Phi=(\phi_{1},\ldots,\phi_{k}),\quad\Psi=(\psi_{1},\ldots,\psi_{k}),\\ \phi_{i}&\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}N_{n}(0_{n},r_{i}W_{\phi}),\quad\psi_{i}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}N_{n}(0_{n},(1-r_{i})I_{n})\;\text{ for $i=1,\ldots,k$,}\end{split}$ (6)
@@
where the regression coefficient matrix $B_{S}$ is $(p+1)\times k$, $G$ is $n\times k$ comprising spatial effects, $M$ is $k\times k$ and invertible that induces correlation between outcomes, $W_{\phi}$ is known $n\times n$ and positive definite capturing spatial autocorrelation, and $R=\mbox{diag}(r_{1},\ldots,r_{k})$ with $r_{i}\in[0,1]$ for $i=1,\ldots k$. The spatial residuals $G$ and the non-spatial residuals $E_{S}$ are modeled in the prior as linear transformations through $M$ of the spatial latent factors $\phi_{1},\ldots,\phi_{k}$ and non-spatial latent factors $\psi_{1},\ldots,\psi_{k}$, respectively, allowing for both spatial and non-spatial heterogeneity. Fixing $M=I_{k}$ results in modeling the outcomes $y_{1},\ldots,y_{k}$ as independent. Similar to the role of $\rho_{i}$ in (3), $r_{i}$ represents the spatial proportion of variance in the $i$th latent process, and setting $R=O$ reduces to the non-spatial likelihood in (4). $W_{\phi}^{-1}$ represents the assumed spatial structure; common constructions include the conditionally autoregressive model (CAR), simultaneous autoregressive model (SAR), or the directed acyclic graph auto-regressive model (DAGAR) developed in Datta et al. (2019). Despite accounting for spatial heterogeneity in the data generation process, (6) is misspecified because the prior on $G$ does not account for the dependency between the confounders $Z$ and covariates $X$.
We specify the prior for (6) as $\pi(B_{S},G,M,R)=N_{nk}(\operatorname{vec}(G)\,|\,0_{nk},M^{\mathrm{\scriptscriptstyle T}}RM\otimes W_{\phi})\times\pi(M)\times\pi(R)$. For $\pi(M)$, we let $\Sigma_{S}=M^{\mathrm{\scriptscriptstyle T}}M$, where $M$ is the upper-triangular Cholesky factor, and set $\pi(\Sigma_{S})=\mbox{IW}(\Sigma_{S}\,|\,v\Sigma_{0},v)$ such that $\mathbb{E}[\Sigma^{-1}_{S}]=\Sigma^{-1}_{0}$. Hence, $\pi(M)\propto\mbox{IW}(M^{\mathrm{\scriptscriptstyle T}}M\,|\,v\Sigma_{0},v)\times\Pi_{i=1}^{k}M_{ii}^{k-i+1}$, where $\Pi_{i=1}^{k}M_{ii}^{k-i+1}$ is proportional to the Jacobian $\left|\frac{\partial\Sigma_{S}}{\partial M_{ij}}\right|$. Following Simpson et al. (2017), we specify a penalized complexity (PC) prior $\pi(R)=\mbox{PC}(R\,|\,\lambda_{R})\propto\lambda_{R}\exp(-\lambda_{R}\sqrt{2\text{KLD}(R)})$ that discourages oversmoothing, where $\lambda_{R}>0$ is a flexibility hyperparameter and $\text{KLD}(R)$ is the Kullback-Leiber divergence of $\mbox{N}_{nk}(0,(I_{k}-R)\otimes I_{n}+R\otimes W_{\phi})$ from $\mbox{N}_{nk}(0,I_{nk})$. Here, $2\text{KLD}(R)=-n\sum_{i=1}^{k}r_{i}+\left(\sum_{i=1}^{k}r_{i}\right)\left(\sum_{j=1}^{n}\lambda_{i}^{-1}\right)-\sum_{i=1}^{k}\sum_{j=1}^{n}\log(1-r_{i}+r_{i}\lambda_{j}^{-1})$, where $\lambda_{1},\ldots,\lambda_{n}$ are the eigenvalues of $W_{\phi}^{-1}$. Small values of $\lambda_{R}$ represent a vague prior while larger values of $\lambda_{R}$ induce heavier shrinkage towards the non-spatial model. A probabilistic statement such as $\mathbb{P}(\sum_{i=1}^{n}r_{i}\leq V)=q$ where $V,q\in(0,1)$ can used as a heuristic to derive $\lambda_{R}$ (Simpson et al., 2017). Samples can be obtained from $\mbox{PC}(R\,|\,\lambda_{R})$ via rejection sampling by repeatedly drawing $r_{1}^{\star},\ldots,r_{k}^{\star}\overset{i.i.d.}{\sim}\mbox{Unif}(0,1)$ and accepting $R_{\star}=\mbox{diag}(r_{1}^{\star},\ldots,r_{k}^{\star})$ with probability $\exp(-\lambda_{R}\sqrt{2\mbox{KLD}(R_{\star}})$. In the simulation and data analysis sections, we use $\lambda_{R}=0.01$, corresponding to a vague prior on $R$. The complete prior is
@@
$\begin{split}\pi(B_{S},G,M,R)&\propto\mbox{N}_{nk}(\operatorname{vec}(G)\,|\,0,M^{\mathrm{\scriptscriptstyle T}}RM\otimes W_{\phi})\\ &\qquad\times\mbox{IW}(M^{\mathrm{\scriptscriptstyle T}}M\,|\,v,v\Sigma_{0})\times\mbox{PC}(R\,|\,\lambda_{R})\times\Pi_{i=1}^{k}M_{ii}^{k-i+1}.\end{split}$ (7)
For tractable analysis, we contrast the posterior profile of $B_{S}$ conditioned on $M$ and $R$ to the posterior of $B_{NS}$ in the non-spatial model of (4) to study the effect of adding the spatial effect $G$ to the analysis model on point estimation and uncertainty quantification.
2.4. Estimation of Fixed Effects
Under the non-spatial analysis model in (4), a natural point estimate for $B$ is $\widehat{B}_{NS}=\mathbb{E}[B_{NS}\,|\,Y,X_{1}]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}Y$, which is equivalent to separate univariate ordinary least squares estimates for each outcome. For the spatial analysis model given by (6), let $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$and $\widehat{G}=\mathbb{E}[G\,|\,Y,M,R,X_{1}]$. Section LABEL:append:spatial_condpost_derivation derives the conditional posterior $\pi(B_{S},\gamma\,|\,Y,M,R,X_{1})$ and the conditional spatial estimator $\widetilde{B}_{S}=\mathbb{E}[B_{S}\,|\,Y,M,R,X_{1}]$ is
$\widetilde{B}_{S}=\widehat{B}_{NS}-(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\widehat{G},\quad\operatorname{vec}({\widehat{G}})=(M^{\mathrm{\scriptscriptstyle T}}\otimes I_{n})L^{-1}\operatorname{vec}\left((I_{n}-H)YM^{-1}\right),$ (8)
and $L=I_{k}\otimes(I_{n}-H)+(R^{-1}(I_{k}-R))\otimes W_{\phi}^{-1}$. The term $(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\widehat{G}$ is the adjustment to the estimation of $B$ from adding a spatial effect that is composed of an inseparable interaction of latent spatial autocorrelation and dependence between results through $W_{\phi}^{-1}$, $M$, and $R$. The spatial effects $G$ absorb spatial information from $Y$ through the residuals $(I_{n}-H)Y$, as $\widetilde{B}_{S}$ can be interpreted as the estimate of $B$ under the non-spatial model in (4) with $Y_{\star}=Y-\mathbb{E}[G\,|\,Y,M,R,X_{1}]$ as the response. The non-spatial estimator can be recovered under the spatial analysis model in (6) by the relation $\widehat{B}_{NS}=\mathbb{E}[B_{S}+(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}G\,|\,Y,M,R,X_{1}]$.
@@ 3. Multivariate Spatial Confounding
3.1. Analysis Model Posterior Variance Inflation
We first compare posterior inference under the Bayesian spatial analysis model of (6) and non-spatial analysis model of (4). Hodges and Reich (2010) argue that unbounded variance inflation occurs in the posterior distribution of the fixed effects as the ratio of non-spatial to spatial variance approaches zero in a Bayesian linear model. To simplify the analysis of spatial smoothness in the multivariate BYM2 model of (6), we let $R=rI_{k}$ such that $r\in(0,1)$ represents the proportion of spatial variation, and define the smoothing ratio as $\frac{1-r}{r}$. Spatial smoothness increases as the smoothing ratio approaches 0, or equivalently, $r\rightarrow 1^{-}$. We show that posterior variance inflation is limited in the multivariate BYM2 model of (6).
By spectral decomposition, there exist an orthogonal matrix $Q\in\mathbb{R}^{n\times n}$ and a diagonal matrix $K=\mbox{diag}(k_{1},\ldots,k_{n})$ with $k_{i}\geq 0$ for all $i=1,\ldots,n$ such that $W_{\phi}^{1/2}(I-H)W_{\phi}^{1/2}=QKQ^{\mathrm{\scriptscriptstyle T}}$. Let $U=W^{1/2}_{\phi}Q$ and $K^{\star}$ be a $n\times n$ diagonal matrix such that $K^{\star}_{ii}=\mbox{I}(k_{i}>0)$ for $i=1,\ldots,n$, where $I(\cdot)$ is the indicator function. The derivations in Section LABEL:append:spatial_condpost_derivation reveal
$\lim_{r\rightarrow 1^{-}}\operatorname{Var}(\operatorname{vec}(B_{S})\,|\,Y,M,R=rI_{k},X_{1})=(M^{\mathrm{\scriptscriptstyle T}}M)\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}W_{\star}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1},$ (9)
where $W_{\star}=W_{\phi}-UK^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Thus, the conditional posterior variance of the fixed effects $B_{S}$ cannot grow infinitely as the residual error becomes predominantly spatial. Similar calculations reveal $\lim_{r\rightarrow 1^{-}}\widetilde{B}_{S}=\widehat{B}_{NS}-(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}UK^{\star}U^{-1}Y$. The limiting between-outcome dependence is encoded by $M^{\mathrm{\scriptscriptstyle T}}M$, another potential source of variance inflation.
However, the limiting distribution of $M^{\mathrm{\scriptscriptstyle T}}M$ is also stable and depends on the structure of $W_{\phi}^{-1}$ through $U$ and $K^{\star}$. Section LABEL:append:limiting_distributions derives the limiting distribution of $M^{\mathrm{\scriptscriptstyle T}}M$ as
$\lim_{r\rightarrow 1^{-}}\pi(M^{\mathrm{\scriptscriptstyle T}}M\,|\,Y,R=rI_{k},X_{1})=\mbox{IW}(M^{\mathrm{\scriptscriptstyle T}}M\,|\,v+n-p-1,v\Sigma_{0}+Y^{\mathrm{\scriptscriptstyle T}}U^{-\mathrm{\scriptscriptstyle T}}K^{\star}U^{-1}Y),$ (10)
@@ 3.1. Analysis Model Posterior Variance Inflation
illuminating the effect of a minimal smoothing ratio on the scale of the conditional variance and between-outcome dependence in the multivariate BYM2 model. The asymptotic results of (9) and (10) indicate that the posterior variance of the fixed effects in $B_{S}$ does not explode when the smoothing ratio is near zero, contrary to the case in Hodges and Reich (2010). Statistical inference remains intact under extreme spatial smoothing, unlike the pathological case of perfect multicollinearity in covariates, although (9) and (10) assume $W_{\phi}^{-1}$ is invertible. This result aligns with Wu and Banerjee (2025) who provide the limiting distribution of the total error variance as the smoothing ratio approaches zero in a univariate BYM2 model.
3.2. Unconditional and Conditional Association Estimands
We now consider the estimators $\widetilde{B}_{S}$ and $\widehat{B}_{NS}$ as point estimators derived from the Bayesian analysis models in Section 2.4 and consider their frequentist properties under the data generation model in (3). Although all parameters in (3) are fixed, there is no universal consensus on the appropriate estimand for assessing the association between the covariates and response. For example, Zimmerman and Ver Hoef (2022) and Reich et al. (2021) consider the association between the covariates and response conditional on the confounder. In our setting, the parameter $B$ represents the association between $Y$ and $X$ conditional on $Z$. The conditional expectations of $\widehat{B}_{NS}$ and $\widetilde{B}_{S}$ under the data generation model in (3) are $\mathbb{E}_{DG}[\widehat{B}_{NS}\,|\,Z,X_{1}]=B+(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}Z$ and $\mathbb{E}_{DG}[\widetilde{B}_{S}\,|\,Z,X_{1}]=B+(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}(Z-\mathbb{E}_{DG}[\widehat{G}\,|\,Z,X_{1}])$, where $\mathbb{E}_{DG}[\operatorname{vec}(\widehat{G})\,|\,Z,X_{1}]=(M^{\mathrm{\scriptscriptstyle T}}\otimes I_{n})L^{-1}\operatorname{vec}((I_{n}-H)ZM^{-1})$. From this perspective, bias already exists in the non-spatial model from collinearity between $X$ and $Z$, and the spatial random effect adjusts the bias based on a product between $X$ and the projection of $Z$ onto the orthogonal complement of $X$ that is weighted according to both spatial structure and between-outcome covariance. The effect of this adjustment on the existing bias depends on the structure of $X$, $Z$, and the analysis parameters $M,R,W_{\phi}^{-1}$.
On the other hand, restricted spatial regression approaches such as those discussed in Hodges and Reich (2010); Hughes and Haran (2013); Hanks et al. (2015) target the unconditional association between the predictors and response, which is represented by $\mathbb{E}_{DG}[Y\,|\,X_{1}]=X(B+D)$ in the multivariate BYM2 model, where $D=(\delta_{0},D_{1}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ represents the linear dependency between the covariates $X_{1}$ and confounder $Z$ that is absorbed into the mean of $Y$. Taking the unconditional expectation of the spatial estimator in (8) under the data generation model in (3) reveals $\mathbb{E}_{DG}[\widehat{B}_{NS}]=\mathbb{E}_{DG}[\widetilde{B}_{S}]=B+D$, so both estimators are unbiased, regardless of the values of $M$, $R$, and $W_{\phi}^{-1}$ used to construct $\widetilde{B}_{S}$.
Although we remain agnostic on this dichotomy, we focus on the estimation properties of $\widehat{B}_{NS}$ and $\widetilde{B}_{S}$ under the data generation model over repeated realizations of $(X_{1},Z,Y)$. In the context of disease mapping, treating $X_{1}$ and $Z$ as stochastic is logical when observed and latent determinants of health, such as population characteristics or environmental factors, may dynamically change when measured alongside the health outcome or are estimated with finite precision. We next turn to whether the spatial analysis model improves estimation precision across multiple datasets of complete health outcome and disease risk maps when compared to the non-spatial analysis model.
3.3. Estimation Over Multiple Datasets
The spatial BYM2 model improves the mean-squared error of point estimates of $B$ over repeated realizations of $(X_{1},Z,Y)$ if it correctly accounts for the variance in the predictors and confounding effects. Analysis in Section LABEL:append:spatial_variance reveals that under the data generation model in (3), if $V_{\phi}=W_{\phi}$ and $\widetilde{B}_{S}=\mathbb{E}[B_{S}\,|\,Y,M=A,R=P,X_{1}]$, then
$\operatorname{Var}_{DG}(\widetilde{B}_{S,ij})<\operatorname{Var}_{DG}(\widehat{B}_{NS,ij})\;\text{and}\;\mathbb{E}_{DG}\left[(\widetilde{B}_{S,ij}-B_{ij})^{2}\right]<\mathbb{E}_{DG}\left[(\widehat{B}_{NS,ij}-B_{ij})^{2}\right]$ (11)